# Characteristic Root Analysis and Regularization for Linear Time Series Forecasting

### v1

@@ 2 Background
Time series forecasting aims to predict future values by learning patterns like trends and seasonality from historical data. Various modeling paradigms have emerged to tackle long-term forecasting challenges (Kong et al., 2025). Transformer-based models have been prominent: Nie et al. (2022) segments series into patch-level tokens with channel-wise modeling, while Zhou et al. (2022) combines seasonal-trend decomposition with transformers. Woo et al. (2023) proposes time-index models that condition on future timestamps to support both interpolation and extrapolation. Beyond time-domain methods, frequency-based approaches like Xu et al. (2023) apply low-pass filtering and complex linear layers to enhance long-term modeling. In parallel, minimalist models have gained traction: Lin et al. (2024) achieves strong results with just 1,000 parameters via structured sparsity, and Zeng et al. (2023) shows that simple linear models can outperform transformers by separately modeling trend and seasonality. Recent evaluations (Brigato et al., 2025) show that no single method dominates across all settings, emphasizing the need for diverse benchmarks and context-aware model selection. Motivated by this, we pursue a theoretical perspective on time series forecasting, aiming for models that balance simplicity, interpretability, and robustness across diverse temporal dynamics.
@@ 3.2 Data Scaling Property Under Noisy Observations
For a linear model forecasting Gaussian white noise, the learned weights converge at a rate proportional to $\mathcal{O}(1/\sqrt{T})$, where $T$ is the length of the observed time series.
@@ 6 Conclusions
This paper presents a comprehensive study of linear models for time series forecasting, focusing on the role of characteristic roots in shaping model expressivity. We propose two complementary methods, Rank Reduction and Root Purge, for improved forecasting through root identification. Experiments validate our theoretical claims and demonstrate the effectiveness of both methods across a range of forecasting tasks. A more detailed discussion and limitations are provided in Appendix F.
@@
Appendix A Summary of Road Map and Notations
A.1 Road Map
Figure 5: Full road map of the paper and its main contributions (full scale).
This diagram summarizes the logical flow of the study, showing how the paper moves from theory to modeling principles and then to algorithm design. It begins with generalization foundations derived from the characteristic roots of linear recurrence relations, which guide the analysis of modeling principles (including lookback window, prediction horizon, instance normalization, channel independence and data scaling laws) under both noise-free and noisy conditions. These theoretical insights motivate the development of two complementary algorithmic strategies—rank reduction (using reduced rank regression or direct weight rank reduction) and the proposed root purge method—which in turn connect back to the characteristic roots, forming a closed loop between theoretical analysis and practical algorithm design. Figure 5 thus provides a roadmap for understanding how the paper integrates classical linear systems theory with modern learning techniques to build robust, interpretable and data-efficient time series forecasting models.
@@ A VAR Perspective on CI
CI (Channel-Independent)
Description All channels share the same linear representation
@@
INC (Individual Channel)
Description Each channel uses a separate linear model
@@
CD (Channel-Dependent)
Description All channels are jointly modeled using a linear model
@@ D.3 Scaling Property on White Noise Series
Let $\mathbf{X}\in\mathbb{R}^{N\times L}$ and $\mathbf{Y}\in\mathbb{R}^{N\times H}$ be random matrices with jointly Gaussian vectorizations:
$\begin{pmatrix}\operatorname{vec}(\mathbf{X})\\ \operatorname{vec}(\mathbf{Y})\end{pmatrix}\sim\mathcal{N}\left(\mathbf{0},\begin{pmatrix}\mathbf{\Sigma}_{\mathbf{X}}&\mathbf{\Sigma_{XY}}\\ \mathbf{\Sigma_{XY}^{\top}}&\mathbf{\Sigma_{Y}}\end{pmatrix}\right)$,
@@
The informal claim of Proposition 1 that the weights learned by a linear model forecasting Gaussian white noise converge at a rate proportional to $\mathcal{O}(1/\sqrt{T})$, where $T$ is the length of the observed time series, can be formally justified by the asymptotic normality theorem for least-squares estimators under joint Gaussian assumptions. In this context, each observation in time can be treated as an independent sample, allowing the time horizon $T$ to play the role of the sample size $N$ in the multivariate analysis. The theorem establishes that, under suitable regularity conditions, the least-squares estimator $\hat{\mathbf{W}}$ converges in distribution to a normal random variable centered at the true coefficient matrix $\mathbf{W}_{\star}$, with convergence rate $\mathcal{O}(1/\sqrt{N})$.
When the time series data consist of Gaussian white noise, the assumptions of the theorem are particularly well-suited. The proposition provides a rigorous statistical foundation for the empirical observation that forecasting models trained on longer time series tend to yield more stable and accurate parameter estimates. However, the convergence rate of the estimator is sublinear—specifically, on the order of $\mathcal{O}(1/\sqrt{T})$—which implies that a disproportionately large amount of data is required to significantly reduce the estimation variance. This highlights an inherent inefficiency in learning from high-variance signals like white noise: although increasing the sample size improves estimation accuracy, the marginal gains diminish. Therefore, to overcome the adverse effects introduced by noise, a substantially large dataset is often necessary.
@@ D.7 Analysis of Root Purge
$J(\mathbf{W})=\left\|\mathbf{Y}-\mathbf{X}\mathbf{W}\right\|_{F}^{2}+\lambda\left\|(\mathbf{Y}-\mathbf{X}\mathbf{W})\mathbf{W}\right\|_{F}^{2},$
admits a special structure. Here, $\mathbf{W}\in\mathbb{R}^{L\times L}$ is now a square matrix.
@@
To further illustrate the relationship between Root Purge and Rank Reduction, we analyze the singular value spectra of $\mathbf{W}^{*}$ and $\mathbf{W}_{\text{OLS}}$. This leads to the following result:
Proposition 14 (Singular Value Shrinkage under Root Purge).
$\sigma_{i}(\mathbf{W}^{*})\leq\sigma_{i}(\mathbf{W}_{\text{OLS}})$
for all $i$, where $\sigma_{i}(\cdot)$ refers to the $i$-th largest singular value of a matrix.
@@
To prove the proposition, we begin by introducing an alternative algorithm—Iteratively Reweighted Least Squares (IRLS)—to solve the following optimization problem:
$J(\mathbf{W})=\|\mathbf{Y}-\mathbf{XW}\|_{F}^{2}+\lambda\|(\mathbf{Y}-\mathbf{XW})\mathbf{W}\|_{F}^{2}.$
  1. 1.
    Initialize: Set the initial weight matrix $\mathbf{W}_{0}=(\mathbf{X}^{\top}\mathbf{X})^{-1}\mathbf{X}^{\top}\mathbf{Y}$;
  2. 2.
    For each iteration $k$:
    1. (a)
      Compute the residual:
      $\mathbf{R}_{k}=\mathbf{Y}-\mathbf{X}\mathbf{W}_{k};$
    2. (b)
      Update the weights:
      $\mathbf{W}_{k+1}=\left(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{R}_{k}^{\top}\mathbf{R}_{k}\right)^{-1}\mathbf{X}^{\top}\mathbf{Y};$
  3. 3.
    Repeat until convergence (e.g., $\|\mathbf{W}_{k+1}-\mathbf{W}_{k}\|_{F}<\epsilon$).
To establish convergence of the algorithm, we define the following surrogate function:
$Q(\mathbf{W};\mathbf{W}_{k})=\|\mathbf{Y}-\mathbf{XW}\|_{F}^{2}+\lambda\cdot\mathrm{tr}(\mathbf{W}^{\top}\mathbf{R}_{k}^{\top}\mathbf{R}_{k}\mathbf{W}).$
Observe that:
  • $Q(\mathbf{W}_{k};\mathbf{W}_{k})=J(\mathbf{W}_{k})$;
  • $Q(\mathbf{W};\mathbf{W}_{k})\geq J(\mathbf{W})$ for all $\mathbf{W}$, since the second term in $Q$ majorizes that in $J$ when linearized around $\mathbf{W}_{k}$.
At each iteration, the IRLS update solves:
$\mathbf{W}_{k+1}=\arg\min_{\mathbf{W}}Q(\mathbf{W};\mathbf{W}_{k}),$
which ensures:
$Q(\mathbf{W}_{k+1};\mathbf{W}_{k})\leq Q(\mathbf{W}_{k};\mathbf{W}_{k})=J(\mathbf{W}_{k}).$
Since $J(\mathbf{W}_{k+1})\leq Q(\mathbf{W}_{k+1};\mathbf{W}_{k})$, it follows that:
$J(\mathbf{W}_{k+1})\leq J(\mathbf{W}_{k}),$
i.e., the objective value decreases monotonically. Because $J(\mathbf{W})\geq 0$, the sequence $\{J(\mathbf{W}_{k})\}$ converges to some limit $J^{*}\geq 0$.
Assume that $\{\mathbf{W}_{k}\}$ is bounded, then it has at least one accumulation point. Let $\mathbf{W}^{*}$ be such a point. By continuity of the update rule, $\mathbf{W}^{*}$ satisfies the fixed-point condition:
$\mathbf{W}^{*}=\left(\mathbf{X}^{\top}\mathbf{X}+\lambda(\mathbf{Y}-\mathbf{XW}^{*})^{\top}(\mathbf{Y}-\mathbf{XW}^{*})\right)^{-1}\mathbf{X}^{\top}\mathbf{Y}.$
To analyze the spectral behavior of $\mathbf{W}^{*}$, define:
$\mathbf{A}=\mathbf{X}^{\top}\mathbf{X},\quad\mathbf{R}^{*}=\mathbf{Y}-\mathbf{X}\mathbf{W}^{*},\quad\mathbf{B}=\lambda\mathbf{R}^{*\top}\mathbf{R}^{*}\succeq 0.$
Then,
$\mathbf{W}^{*}=(\mathbf{A}+\mathbf{B})^{-1}\mathbf{X}^{\top}\mathbf{Y},\quad\mathbf{W}_{\text{OLS}}=\mathbf{A}^{-1}\mathbf{X}^{\top}\mathbf{Y}.$
Since $\mathbf{A}+\mathbf{B}\succeq\mathbf{A}$, we have $(\mathbf{A}+\mathbf{B})^{-1}\preceq\mathbf{A}^{-1}$. By Weyl’s monotonicity theorem for singular values of matrix products with positive definite matrices, it follows that:
$\sigma_{i}(\mathbf{W}^{*})\leq\sigma_{i}(\mathbf{W}_{\text{OLS}}),\quad\forall\ i.$
@@
Furthermore, we can see that larger values of $\lambda$ lead to a stronger suppression of the singular values of $\mathbf{W}^{*}$. This theoretical behavior is clearly reflected in our empirical results, as shown in Figure  3 and Figure 13, where increasing $\lambda$ progressively attenuates smaller singular values. This supports the interpretation that Root Purge acts similarly to rank reduction techniques by selectively suppressing spurious or noisy components (roots) in the solution. Hence, Root Purge can be seen as a principled mechanism for regularizing overparameterized models by promoting low-rank structure.
@@
The gradient $\nabla J(\mathbf{W})$ is Lipschitz continuous with constant $L$ satisfying [^8] The Lipschitz constant $L$ for $\nabla J(\mathbf{W})$ can be estimated as two parts: root-seeking term $\|\mathbf{Y}-\mathbf{XW}\|_{F}^{2}$ contributes to $2\sigma_{\max}^{2}(\mathbf{X})$, and root-purging term $\|(\mathbf{Y}-\mathbf{XW})\mathbf{W}\|_{F}^{2}$ yields the $\lambda$-scaled component $2\lambda\left[\|\mathbf{X}\|_{2}(\|\mathbf{Y}\|_{2}+\|\mathbf{X}\|_{2}M)M+(\|\mathbf{Y}\|_{2}+\|\mathbf{X}\|_{2}M)^{2}\right]$ via triangle inequalities and submultiplicativity.:
$L\leq 2\sigma_{\text{max}}^{2}(\mathbf{X})+2\lambda(2M^{2}\|\mathbf{X}\|_{2}^{2}+3M\|\mathbf{X}\|_{2}\|\mathbf{Y}\|_{2}+\|\mathbf{Y}\|_{2}^{2})$ (11)
@@
  • $\sigma_{\text{max}}(\mathbf{X})$ is the largest singular value of $\mathbf{X}$;

### v2

@@ 2 Background
Time series forecasting aims to predict future values by learning patterns like trends and seasonality from historical data. Various modeling paradigms have emerged to tackle long-term forecasting challenges (Kong et al., 2025). Transformer-based models have been prominent: Nie et al. (2022) segments series into patch-level tokens with channel-wise modeling, while Zhou et al. (2022) combines seasonal-trend decomposition with transformers. Woo et al. (2023) proposes time-index models that condition on future timestamps to support both interpolation and extrapolation. Beyond time-domain methods, frequency-based approaches like Xu et al. (2023) apply low-pass filtering and complex linear layers to enhance long-term modeling. In parallel, minimalist models have gained traction: Lin et al. (2024) achieves strong results with just 1,000 parameters via structured sparsity, and Zeng et al. (2023) shows that simple linear models can outperform transformers by separately modeling trend and seasonality. Recent evaluations (Brigato et al., 2025) show that no single method dominates across all settings, emphasizing the need for diverse benchmarks and context-aware model selection. Motivated by this, we pursue a theoretical perspective on time series forecasting, aiming for models that balance simplicity, interpretability, and robustness across diverse temporal dynamics. To see a detailed related work and further background, please refer to Appendix B.1.
@@ 3.2 Data Scaling Property Under Noisy Observations
For a linear model forecasting a zero-mean noise with finite second moments, the learned weights converge at a rate proportional to $\mathcal{O}(1/\sqrt{T})$, where $T$ is the length of the observed time series.
@@ 6 Conclusions
This paper presents a comprehensive study of linear models for time series forecasting, focusing on the role of characteristic roots in shaping model expressivity. We propose two complementary methods, Rank Reduction and Root Purge, for improved forecasting through root identification. Experiments validate our theoretical claims and demonstrate the effectiveness of both methods across a range of forecasting tasks. More detailed discussions and limitations are provided in Appendix F.
@@
Appendix A Mind Map, Notations, and Summary of Theorems
A.1 Mind Map
Figure 5: A mind map of the paper describing its logic flow and how we organize this paper.
This mind map summarizes the logical flow of the study, showing how the paper moves from theory to modeling principles and then to algorithm design. It begins with generalization foundations derived from the characteristic roots of linear recurrence relations, which guide the analysis of modeling principles (including lookback window, prediction horizon, instance normalization, channel independence, and data scaling laws) under both noise-free and noisy conditions. These theoretical insights motivate the development of two complementary algorithmic strategies—rank reduction (using reduced rank regression or direct weight rank reduction) and the proposed root purge method—which in turn connect back to the characteristic roots, forming a closed loop between theoretical analysis and practical algorithm design. Figure 5 thus provides a roadmap for understanding how the paper integrates classical linear systems theory with modern learning techniques to build robust, interpretable, and data-efficient time series forecasting models.
@@ A.2 Notation
A.3 Summary of Theorems
In our paper, we had proposed multiple theorems, facts, propositions, and corollaries about linear forecasting models and our proposed methods. To provide a comprehensive idea of how each of them contributes to our analysis and what they mean in an intuitive sense, we summarize them in Table LABEL:appx:tab:theory-summary shown below. For the detailed assumption and proofs, please refer to Appendix D.
Table 4: Summary of Key Theorems, Facts, Propositions, and Corollaries
Theorem Plain-Language Explanation Role and Significance in the Paper
Fact 1 A linear model can accurately represent a time series if the time series’ characteristic roots are a subset of the model’s characteristic roots. 1. Defines core generalization logic: adaptation depends on root subset relationships. 2. Provides theoretical basis for fitting diverse time series.
Fact 2 For linear models with longer forecasting horizons or extended lookback windows, their set of characteristic roots always includes the key roots of the system. 1. Justifies independent modeling per forecast horizon. 2. Shows parameter redundancy enhances flexibility without harming core dynamics.
Proposition 1
(Informal version of Prop. 6)
For a zero-mean noise with finite second moments, convergence rate of OLS weights is $O(1/\sqrt{T})$. 1. Reveals limitations of least-squares in noisy data. 2. Motivates structural regularization (RRR, Root Purge).
Proposition 2
(Informal version of Prop. 7)
Constraining weight matrix $W$ to low-rank is equivalent to projecting input and output into a shared low-dimensional subspace. 1. Basis for RRR/DWRR methods. 2. Explains denoising via low-rank constraints.
Proposition 3 Multi-channel time series with diagonal recurrence per channel can be transformed to a shared-parameter recurrence. 1. Proves effectiveness of channel-independent modeling. 2. Shared weights improve robustness and reduce parameters.
Proposition 4 Two exponential combinations with different characteristic roots that match at $L$ points must match everywhere. 1. Strengthens Fact 1: root subset relation determines expressivity. 2. Provides generalization guarantee.
Proposition 5 Characteristic polynomial of minimum-order recurrence divides higher-order recurrences. 1. Deepens Fact 2 from polynomial theory. 2. High-order models only add redundant roots.
Proposition 6
(Formal version of Prop. 1)
For a zero-mean noise with finite second moments, convergence rate of OLS weights is $O(1/\sqrt{T})$. 1. Reveals limitations of least-squares in noisy data. 2. Motivates structural regularization (RRR, Root Purge).
Proposition 7
(Formal version of Prop. 2)
Constraining weight matrix $W$ to low-rank is equivalent to projecting input and output into a shared low-dimensional subspace. 1. Basis for RRR/DWRR methods. 2. Explains denoising via low-rank constraints.
Proposition 8 DWRR’s low-rank approximation of OLS is optimal (Frobenius norm), error = sum of discarded singular values squared. 1. Proves DWRR optimality. 2. Guides engineering implementation.
Proposition 9 RRR error is bounded by minimal singular values of input and output data. 1. Quantifies RRR approximation accuracy. 2. Guides rank selection.
Proposition 10 If true matrix is low-rank, truncated SVD error after noise is controlled by noise spectral norm. 1. Provides matrix denoising interpretation. 2. Low-rank constraints robustly isolate noise.
Proposition 11 Error of estimating true principal subspace decreases with sample size. 1. Ensures statistical reliability of rank reduction. 2. Supports tuning rank $\rho$ with validation data.
Proposition 12 OLS is a critical point of Root Purge objective iff residual and weights satisfy null-space condition. 1. Shows relationship between Root Purge and OLS. 2. Guides regularization for noise suppression.
Proposition 13 Deviation between Root Purge and OLS weights is bounded by minimal singular values and regularization $\lambda$. 1. Quantifies correction magnitude. 2. Ensures stability of Root Purge.
Proposition 14 Each singular value of Root Purge weights $\leq$ corresponding OLS weight singular value. 1. Shows Root Purge implicitly enforces low-rankness. 2. Connects Root Purge to rank reduction.
Proposition 15 Gradient descent with proper step size converges to a stationary point of Root Purge objective. 1. Ensures feasibility of Root Purge optimization. 2. Justifies using OLS warm starts.
Proposition 16 If weight matrix approximation error = $\varepsilon$, root perturbation = $O(\varepsilon)$. 1. Connects weight approximation to root stability. 2. Ensures regularized model preserves core dynamics.
Corollary 1 Sinusoids, polynomial, exponential, and combinations can be represented by linear difference equations. 1. Shows linear models fit common temporal patterns. 2. Supports long-term forecasting.
Corollary 2 A linear model represents a time series iff the time series’ PV extension is a subextension of the model’s PV extension. 1. Uses differential Galois theory to characterize model expressivity. 2. Shows linear models’ generality for complex dynamics.
@@ A VAR Perspective on CI
CI (Channel-Independent) Description All channels share the same linear representation
@@
INC (Individual Channel) Description Each channel uses a separate linear model
@@
CD (Channel-Dependent) Description All channels are jointly modeled using a linear model
@@ D.3 Scaling Property on White Noise Series
Let $\mathbf{X}\in\mathbb{R}^{N\times L}$ and $\mathbf{Y}\in\mathbb{R}^{N\times H}$ be random matrices with the following moment conditions:
$\mathbb{E}\left[\begin{pmatrix}\operatorname{vec}(\mathbf{X})\\ \operatorname{vec}(\mathbf{Y})\end{pmatrix}\right]=\mathbf{0},\quad\mathbb{E}\left[\begin{pmatrix}\operatorname{vec}(\mathbf{X})\\ \operatorname{vec}(\mathbf{Y})\end{pmatrix}\begin{pmatrix}\operatorname{vec}(\mathbf{X})\\ \operatorname{vec}(\mathbf{Y})\end{pmatrix}^{\top}\right]=\begin{pmatrix}\bm{\Sigma}_{\mathbf{X}}&\bm{\Sigma}_{\mathbf{XY}}\\ \bm{\Sigma}_{\mathbf{XY}}^{\top}&\bm{\Sigma}_{\mathbf{Y}}\end{pmatrix}$,
@@
The informal claim of Proposition 1 that the weights learned by a linear model forecasting a zero-mean noise with finite second moments converge at a rate proportional to $\mathcal{O}(1/\sqrt{T})$, where $T$ is the length of the observed time series, can be formally justified by the asymptotic normality theorem for least-squares estimators. In this context, each observation in time can be treated as an independent sample, allowing the time horizon $T$ to play the role of the sample size $N$ in the multivariate analysis. The theorem establishes that, under suitable regularity conditions, the least-squares estimator $\hat{\mathbf{W}}$ converges in distribution to a normal random variable centered at the true coefficient matrix $\mathbf{W}_{\star}$, with convergence rate $\mathcal{O}(1/\sqrt{N})$.
When the time series data consist of a zero-mean noise with finite second moments, the assumptions of the theorem are particularly well-suited. The proposition provides a rigorous statistical foundation for the empirical observation that forecasting models trained on longer time series tend to yield more stable and accurate parameter estimates. However, the convergence rate of the estimator is sublinear—specifically, on the order of $\mathcal{O}(1/\sqrt{T})$—which implies that a disproportionately large amount of data is required to significantly reduce the estimation variance. This highlights an inherent inefficiency in learning from high-variance signals like white noise: although increasing the sample size improves estimation accuracy, the marginal gains diminish. Therefore, to overcome the adverse effects introduced by noise, a substantially large dataset is often necessary.
@@ D.7 Analysis of Root Purge
$J(\mathbf{W})=\left\|\mathbf{Y}-\mathbf{X}\mathbf{W}\right\|_{F}^{2}+\lambda\left\|(\mathbf{Y}-\mathbf{X}\mathbf{W})\mathbf{W}\right\|_{F}^{2}=\left\|\mathbf{Y}-\mathbf{X}\mathbf{W}\right\|_{F}^{2}+\lambda\left\|\mathbf{R}\mathbf{W}\right\|_{F}^{2},$
admits a special structure. Here, $\mathbf{W}\in\mathbb{R}^{L\times L}$ is now a square matrix. In practice, the optimization of $J(\mathbf{W})$ can also be carried out using a stop-gradient strategy, where the residual in the root purge term $\mathbf{R}=\mathbf{Y}-\mathbf{X}\mathbf{W}$ is treated as constant when updating $\mathbf{W}$. Since the stop-gradient case tends to be easier, we mainly analyze the full-gradient version of root purge and give conclusions of stop-gradient version in the remarks.
@@
Remark 5 (Equivalence Condition of Stop-Gradient Root Purge).
For the stop-gradient objective
$\widetilde{J}(\mathbf{W})=\|\mathbf{Y}-\mathbf{XW}\|_{F}^{2}+\lambda\|\mathbf{R}\mathbf{W}\|_{F}^{2},$
where the residual matrix $\mathbf{R}$ is treated as constant during optimization, the same conclusion from Proposition 12 holds.
Remark 6 (Bounded Deviation of Stop-Gradient Root Purge).
For the stop-gradient objective
$\widetilde{J}(\mathbf{W})=\|\mathbf{Y}-\mathbf{XW}\|_{F}^{2}+\lambda\|\mathbf{R}\mathbf{W}\|_{F}^{2},$
where the residual matrix $\mathbf{R}$ is treated as constant during optimization, the gradient term become $\nabla\widetilde{J}(\mathbf{W})=-2\mathbf{X}^{\top}\mathbf{R}+2\lambda\mathbf{R}^{\top}\mathbf{R}\mathbf{W},$ and the bound as stated in Proposition 13 becomes
$\|\mathbf{W}^{*}-\mathbf{W}_{\text{OLS}}\|_{F}\leq\frac{\|\mathbf{E}^{\top}\mathbf{E}\mathbf{W}_{\text{OLS}}\|_{F}}{1/\lambda\cdot\sigma_{\min}(\mathbf{X}^{\top}\mathbf{X})+\sigma_{\min}(\mathbf{E}^{\top}\mathbf{E})}.$
To further illustrate the relationship between Root Purge and Rank Reduction, we have the following results:
Proposition 14 (Anisotropic Shrinkage and Conditional Singular Value Shrinkage at the Stationary Point).
Assume that $\mathbf{X}^{\top}\mathbf{X}\succ 0$, and $\mathbf{W}^{*}$ satisfies the orthogonality condition
$\mathbf{X}^{\top}(\mathbf{Y}-\mathbf{X}\mathbf{W}^{*})\mathbf{W}^{*}=\mathbf{0}.$
Then there exists a symmetric positive semidefinite matrix $\mathbf{S}\preceq\mathbf{I}$ such that
$\mathbf{W}^{*}=\mathbf{S}\,\mathbf{W}_{\mathrm{OLS}},$
where
$\mathbf{S}=\left(\mathbf{I}+(\mathbf{X}^{\top}\mathbf{X})^{-1}\lambda\mathbf{R}^{*\top}\mathbf{R}^{*}\right)^{-1},\qquad\mathbf{R}^{*}=\mathbf{Y}-\mathbf{X}\mathbf{W}^{*}.$
If we further assume that $\mathbf{X}^{\top}\mathbf{X}$ and $\mathbf{R}^{\top}\mathbf{R}$ are simultaneously diagonalizable. Then the unique minimizer $\mathbf{W}^{*}$ satisfies
$\sigma_{i}(\mathbf{W}^{*})\leq\sigma_{i}(\mathbf{W}_{\mathrm{OLS}}),\quad\forall i.$
@@
Let $\mathbf{R}=\mathbf{Y}-\mathbf{X}\mathbf{W}$. Using standard matrix calculus, the gradient of $J$ is given by
$\nabla J(\mathbf{W})=-2\mathbf{X}^{\top}\mathbf{R}+2\lambda\left(\mathbf{R}^{\top}\mathbf{R}\mathbf{W}-\mathbf{X}^{\top}\mathbf{R}\mathbf{W}\mathbf{W}^{\top}\right).$
At a stationary point $\mathbf{W}^{*}$, $\nabla J(\mathbf{W}^{*})=\mathbf{0}$, which yields
$\mathbf{X}^{\top}\mathbf{R}^{*}+\lambda\mathbf{X}^{\top}\mathbf{R}^{*}\mathbf{W}^{*}\mathbf{W}^{*\top}=\lambda\mathbf{R}^{*\top}\mathbf{R}^{*}\mathbf{W}^{*}.$
Under the assumed orthogonality condition $\mathbf{X}^{\top}\mathbf{R}^{*}\mathbf{W}^{*}=\mathbf{0}$, the second term on the left-hand side vanishes, and the above equation reduces to
$(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{R}^{*\top}\mathbf{R}^{*})\mathbf{W}^{*}=\mathbf{X}^{\top}\mathbf{Y}.$
Let $\mathbf{A}=\mathbf{X}^{\top}\mathbf{X}$, $\mathbf{B}=\lambda\mathbf{R}^{*\top}\mathbf{R}^{*}\succeq 0$, and $\mathbf{C}=\mathbf{X}^{\top}\mathbf{Y}$. Then
$\mathbf{W}^{*}=(\mathbf{A}+\mathbf{B})^{-1}\mathbf{C},\qquad\mathbf{W}_{\mathrm{OLS}}=\mathbf{A}^{-1}\mathbf{C}.$
Using the identity
$(\mathbf{A}+\mathbf{B})^{-1}=(\mathbf{I}+\mathbf{A}^{-1}\mathbf{B})^{-1}\mathbf{A}^{-1},$
we obtain
$\mathbf{W}^{*}=\underbrace{(\mathbf{I}+\mathbf{A}^{-1}\mathbf{B})^{-1}}_{\mathbf{S}}\mathbf{W}_{\mathrm{OLS}}.$
Since $\mathbf{A}^{-1}\mathbf{B}\succeq 0$, it follows that $\mathbf{0}\preceq\mathbf{S}\preceq\mathbf{I}$, completing the proof of our first conclusion. Additionally, by further assumptions, $\mathbf{A}$ and $\mathbf{B}$ are simultaneously diagonalizable. Hence there exists an orthogonal matrix $\mathbf{U}$ such that
$\mathbf{A}=\mathbf{U}\operatorname{diag}(a_{i})\mathbf{U}^{\top},\qquad\mathbf{B}=\mathbf{U}\operatorname{diag}(b_{i})\mathbf{U}^{\top},\quad b_{i}\geq 0.$
Therefore,
$(\mathbf{A}+\mathbf{B})^{-1}=\mathbf{U}\operatorname{diag}\!\left(\frac{1}{a_{i}+b_{i}}\right)\mathbf{U}^{\top},\qquad\mathbf{A}^{-1}=\mathbf{U}\operatorname{diag}\!\left(\frac{1}{a_{i}}\right)\mathbf{U}^{\top}.$
Let $\widetilde{\mathbf{C}}=\mathbf{U}^{\top}\mathbf{C}$. Then
$\mathbf{W}^{*}=\mathbf{U}\operatorname{diag}\!\left(\frac{1}{a_{i}+b_{i}}\right)\widetilde{\mathbf{C}},\qquad\mathbf{W}_{\mathrm{OLS}}=\mathbf{U}\operatorname{diag}\!\left(\frac{1}{a_{i}}\right)\widetilde{\mathbf{C}}$.
Since $a_{i}+b_{i}\geq a_{i}$ for all $i$, it follows that
$\frac{1}{a_{i}+b_{i}}\leq\frac{1}{a_{i}},\quad\forall i.$
Consequently, each singular value of $\mathbf{W}^{*}$ is bounded above by the corresponding singular value of $\mathbf{W}_{\mathrm{OLS}}$, which completes the proof. ∎
The first part of Proposition 14 establishes that, under a mild orthogonality condition, the stationary solution $\mathbf{W}^{*}$ can be expressed as a linear transformation of the OLS estimator via a contraction operator $\mathbf{S}\preceq\mathbf{I}$. This result reveals an anisotropic shrinkage mechanism, where different directions in parameter space are suppressed to different degrees depending on the residual covariance. With further assumption in the second part, we can additionally perform a component-wise comparison of singular values between $\mathbf{W}^{*}$ and $\mathbf{W}_{\mathrm{OLS}}$.
@@
The simultaneous diagonalizability of $\mathbf{X}^{\top}\mathbf{X}$ and $\mathbf{R}^{\top}\mathbf{R}$ should be interpreted as an analytical choice of coordinate system rather than a restrictive modeling assumption. Indeed, the result characterizes shrinkage behavior in the common eigenbasis of the data covariance and the residual covariance. In particular, directions associated with large residual energy experience stronger suppression, while directions well explained by the data are preserved. As the regularization parameter $\lambda$ increases, this selective attenuation progressively eliminates low-energy singular modes, effectively reducing the rank of the solution. This theoretical behavior is clearly reflected in our empirical results, as shown in Figure  3 and Figure 14, where increasing $\lambda$ progressively attenuates smaller singular values. This supports the interpretation that Root Purge acts similarly to rank reduction techniques by selectively suppressing spurious or noisy components (roots) in the solution. Hence, Root Purge can be seen as a principled mechanism for regularizing overparameterized models by promoting low-rank structure.
Remark 7 (Anisotropic vs. isotropic shrinkage).
Unlike ridge regression, which applies uniform (isotropic) shrinkage across all directions, the shrinkage operator $\mathbf{S}$ depends on the residual covariance $\mathbf{R}^{*\top}\mathbf{R}^{*}$ and therefore induces a data-dependent, directionally selective contraction of the OLS estimator.
Remark 8 (Norm contraction).
As a direct consequence of $\mathbf{0}\preceq\mathbf{S}\preceq\mathbf{I}$, we have
$\|\mathbf{W}^{*}\|_{2}\leq\|\mathbf{W}_{\mathrm{OLS}}\|_{2},\qquad\|\mathbf{W}^{*}\|_{F}\leq\|\mathbf{W}_{\mathrm{OLS}}\|_{F}.$
Remark 9 (Shrinkage of Stop-Gradient Root Purge).
Consider the stop-gradient objective
$\widetilde{J}(\mathbf{W})=\|\mathbf{Y}-\mathbf{XW}\|_{F}^{2}+\lambda\|\mathbf{R}\mathbf{W}\|_{F}^{2},$
where the residual matrix $\mathbf{R}$ is treated as constant during optimization. We can drop the orthogonality condition that $\mathbf{X}^{\top}(\mathbf{Y}-\mathbf{X}\mathbf{W}^{*})\mathbf{W}^{*}=\mathbf{0}$. We still arrive at the conclusion of Proposition 14 given the rest of assumptions are satisfied.
@@
The gradient $\nabla J(\mathbf{W})$ is Lipschitz continuous with constant $L$ satisfying [^8] The Lipschitz constant $L$ for $\nabla J(\mathbf{W})$ can be estimated as two parts: root-seeking term $\|\mathbf{Y}-\mathbf{XW}\|_{F}^{2}$ contributes to $2\|\mathbf{X}\|_{2}^{2}$, and root-purging term $\|(\mathbf{Y}-\mathbf{XW})\mathbf{W}\|_{F}^{2}$ has two gradient components: $2\lambda\mathbf{R}^{\top}\mathbf{R}\mathbf{W}$ and $2\lambda\mathbf{X}^{\top}\mathbf{R}\mathbf{W}\mathbf{W}^{\top}$. The bound of $L$ can be found by looking at each component separately. The first component give us the bound of $2\lambda\left[2\|\mathbf{X}\|_{2}(\|\mathbf{Y}\|_{2}+\|\mathbf{X}\|_{2}M)M+(\|\mathbf{Y}\|_{2}+\|\mathbf{X}\|_{2}M)^{2}\right]$ and the later give us the bound of $2\lambda\left[2\|\mathbf{X}\|_{2}(\|\mathbf{Y}\|_{2}+\|\mathbf{X}\|_{2}M)M+\|\mathbf{X}\|^{2}_{2}M^{2}\right]$ via triangle inequalities and submultiplicativity.:
$L\leq 2\|\mathbf{X}\|_{2}^{2}+2\lambda(6M^{2}\|\mathbf{X}\|_{2}^{2}+6M\|\mathbf{X}\|_{2}\|\mathbf{Y}\|_{2}+\|\mathbf{Y}\|_{2}^{2})$ (11)